# 3D Hierarchical Panoptic Segmentation in Real Orchard Environments Across Different Sensors

### v1

@@ Abstract
Crop yield estimation is a relevant problem in agriculture, because an accurate crop yield estimate can support farmers’ decisions on harvesting or precision intervention. Robots can help to automate this process. To do so, they need to be able to perceive the surrounding environment to identify target objects. In this paper, we introduce a novel approach to address the problem of hierarchical panoptic segmentation of apple orchards on 3D data from different sensors. Our approach is able to simultaneously provide semantic segmentation, instance segmentation of trunks and fruits, and instance segmentation of plants (a single trunk with its fruits). This allows us to identify relevant information such as individual plants, fruits, and trunks, and capture the relationship among them, such as precisely estimate the number of fruits associated to each tree in an orchard. Additionally, to efficiently evaluate our approach for hierarchical panoptic segmentation, we provide a dataset designed specifically for this task. Our dataset is recorded in Bonn in a real apple orchard with a variety of sensors, spanning from a terrestrial laser scanner to a RGB-D camera mounted on different robotic platforms. The experiments show that our approach surpasses state-of-the-art approaches in 3D panoptic segmentation in the agricultural domain, while also providing full hierarchical panoptic segmentation. Our dataset has been made publicly available at https://www.ipb.uni-bonn.de/data/hops/. We will provide the open-source implementation of our approach and public competiton for hierarchical panoptic segmentation on the hidden test sets upon paper acceptance.
@@ I Introduction
Crop yield estimation is an important task in agriculture, as an accurate crop yield estimate can help farmers making management decisions to increase crop production and optimize key factors such as harvest time and fertilizer use [17, 41, 44], as well as enable precision interventions [2]. Robots can automatize or support many of these interventions, but to do so reliably and robustly, they need to understand their surroundings through the interpretation of sensor data. Tasks such as segmenting, counting, and localizing fruits in orchards are crucial in horticulture, as they are required for replacing human fruit picking, which is an extremely labor-intensive process [3].
@@
Our experiments suggest that (i) our approach can jointly perform semantic, tree instance, and standard instance segmentation on real-world 3D data acquired with different sensors; and (ii) our skip connection scheme allows the CNN to exploit the underlying hierarchical relationship between the individual segmentation tasks, leading to better segmentation performance. Our claims are backed up by the paper and our experimental evaluation. The dataset is publicly available. The implementation of our approach and the dataset competitions will be available upon acceptance.
@@ II Related Work
Panoptic segmentation [15] unifies semantic and instance segmentation provides semantics of background classes, so-called “stuff”, and object-level instances of so-called “things” at the same time. While often images are used for panoptic segmentation in the agricultural domain [43], 3D point clouds generated by RGB-D cameras, LiDARs, or photogrammetric structure-from-motion approaches received increasing research interest [46] due to the capability to extract precise geometric information required for high-precision fruit grasping [22, 16], plant segmentation [30], and fine-grained trait estimation [26]. 3D panoptic segmentation has been actively investigated in the domain of autonomous driving [37, 28, 9, 45], where most approaches use dedicated branches for semantic and instance segmentation using features of a shared encoder. For instance segmentation, these approaches employ either (i) a bottom-up paradigm using instance-wise offset vectors [28, 45] or embeddings [9] for clustering of instances or (ii) a top-down paradigm with object detectors identifying instances that are then grouped via the semantic prediction [37]. For plant phenotyping, we can exploit the hierarchical structure of plants that can be decomposed in individual parts [12], such as plant, leaf and even more fine-grained leaf structure. This hierarchical structure has been exploited in the literature, leveraging skip connections between decoders for hierarchical bottom-up instance prediction tasks using images [34], which is the foundation of our approach to 3D panoptic segmentation.
In contrast to prior approaches for 3D panoptic segmentation, we also explicitly consider the hierarchical structure of the prediction task, which allows us to predict a hierarchical semantic scene interpretation consisting of plant instances as well as fruit instances that can be used for plant monitoring and yield estimation in the agricultural domain.
While most agricultural datasets for semantic interpretation provide only RGB images [21], recently 3D datasets using various sensors became available. In particular, BUP20 [38] provides an RGB-D dataset for instance segmentation of sweet peppers in a glasshouse [39]. For 3D plant phenotyping in real agricultural fields, BonnBeetClouds [25] provides annotated structure-from-motion point clouds consisting of plant and leaf instance of sugar beet plants. Other datasets provide organ-level annotations of plants using point clouds acquired with high-precision LiDAR scanners [36] or x-ray imaging [7] in the lab. The Crops3D dataset [46] provides organ-level annotations of different crop varieties for point cloud data acquired with a TLS in the field and single plant point clouds using structure-from-motion of RGB images or using structured light cameras.
@@ III-B Skip Connections
Skip connections are crucial to ensure feature reusability and address the gradient degradation problem of CNNs. They skip one or more layers, providing a direct gradient flow from the late stages of the decoder to the early stages of the encoder, which is usually hindered by the downsampling operation of the encoder. Ronneberger et al. [35] first proposed the use of encoder-decoder skip connections, where higher-resolutions maps of the encoder are concatenated with the feature maps of the decoder. We aim to extend our previous work on hierarchical panoptic segmentation on RGB images [34] to the 3D point cloud domain, and thus we propose to adopt a hierarchical skip connections scheme to address the underlying hierarchy of segmentation tasks. We propose to connect different decoders directly as shown in Fig. 2, instead of encoder and decoders only.
@@
This skip connection scheme allows us to exploit the hierarchy among tasks, enriching the features propagated to each decoder. In Sec. V-C, we provide extensive experiments on different skip connection schemes to show that this design choice yields superior performance.
@@ IV Our Dataset for
Our dataset, called HOPS, is composed of point clouds collected with three different sensors at Campus Klein-Altendorf near Bonn, Germany. First, we collected point clouds using a TLS placed at multiple locations in the orchard row. Our training and validation set entirely contains point clouds collected with this sensor. Furthermore, we use a few TLS point clouds in the test set, collected in a different year and orchard location compared to the training and validation set. Second, we use point clouds obtained with a bundle adjustment procedure [40] using as input images collected with a UAV equipped with a PhaseOne iXM-100 camera. We flew three missions on the same orchard with a camera angle of 45∘, 90∘, and 135∘ from the ground plane at a height of approximately 20 m. This setup allows us to obtain good coverage of the trees including the lower part of the canopy. The photogrammetric point clouds obtained in this way from the UAV are only included in the test set. Third, we use a RealSense d435i mounted on a mobile robot driving through the orchard rows to collect another part of our test set. Again, to obtain photogrammetric point clouds we use a bundle adjustment pipeline. Furthermore, we label a few point clouds from the Fuji-SfM dataset [10, 11] to obtain a fourth test set from a different setup. This dataset is collected in Agramunt in Catalunia, Spain with a EOS 60D DSLR Canon camera, followed by a bundle adjustment procedure to generate point clouds. In contrast to existing datasets in the agricultural domain [13] [33] [42], we specifically designed HOPS to have a remarkable domain shift between train and test sets, to more closely resemble real-world deployments. For this reason, we only included high-quality point clouds recorded with the TLS in the training set, while the test set includes a variety of sensors.
@@
To label the data, we split the aggregated point cloud of each orchard row in individual tiles with an average of 1 million points each. We manually label each tile in three stages using the online tool provided by segments.ai. We start with one annotator labeling for semantic segmentation and the two level instance segmentation, ensuring consistency among the two instance segmentation labels. Lastly, a different annotator verifies the label quality before moving on to the next tile. Compared to image-based annotations, the labeling process of point clouds is more complex, since an annotator often needs to change the viewpoint to label parts of the point cloud. This results in an average labeling time of 4 hours per tile, with an overall effort consisting of for the labeling and for the verification.
@@ V-A Experimental Setup
Metrics. For semantic segmentation, we compute the intersection over union (IoU) [8] over all five classes of our dataset, which we discussed in Sec. IV, and report the mean IoU (mIoU) in the following tables. For the tree instance segmentation, we compute the panoptic quality over the “tree” class, that we report as PQ. For the standard instance segmentation, we compute the panoptic quality over all classes, and report the mean panoptic quality as PQ in the tables. We also report the overall mean panoptic quality (mPQ) as the average between PQ and PQ.
@@
| Approach | TLS | |||
| mIoU | PQ | PQ | mPQ | |
| MaskPLS [24] | 28.2 | 42.8 | - | 45.6 |
| MaskPLS [24] | - | - | 48.9 | |
| ForestPS [23] | 58.9 | 62.2 | - | 32.3 |
| ForestPS [23] | - | - | 2.3 | |
| Ours | 70.9 | 60.8 | 52.1 | 56.5 |
Training details and parameters. In all experiments, we use AdamW [18] with weight decay of with an initial learning rate of . We set the weights of the loss function to . We train our approach for 500 epochs. We set batch size of 1. We use voxel downsampling to . For training, we use several data augmentation techniques to avoid overfitting on a specific sensor. In particular, we use scale, rotation around all axes, shear, color jittering, and elastic deformation [4]. We tuned all hyperparameters on the validation set.
@@ V-B Experiments on Hierarchical Panoptic Segmentation
The first experiment evaluates the performance of our approach and its outcomes support the claim that our approach can jointly perform semantic, tree instance, and standard instance segmentation on real-world 3D data. For this experiment, we compare our approach to existing baselines for panoptic segmentation: MaskPLS [24] is a transformer-based approach that extends Mask2Former [5] to 3D point clouds, while ForestPS [23] is a convolutional neural network approach based on MinkUNet [6], originally designed for tree segmentation in a forest environment. The baselines are trained for only one of the two instance segmentation tasks at a time, as they do not support hierarchical multi-level instance segmentation. As shown in Tab. II for the test set, and Tab. III for the validation set, our approach outperforms the baselines on both splits in terms of mIoU and on most PQ, while ForestPS achieves the best PQ on most splits. However, our approach performs better than the baselines on mPQ. While our PQ is inferior to ForestPS, our approach also yields good performance on tree instance segmentation, on which the dedicated training of ForestPS fails completely, achieving results below on four out of five splits.
We show qualitative results in Fig. 4. As explained in Sec. IV, only the validation set is recorded in the same time period and with the same sensor as the training set. The test sets, in contrast, are recorded either 2 years later (TLS), or with different sensors (UGV and UAV), or belong to entirely different datasets (SfM). This motivates the performance gap between the validation and the test set. We believe that this aspect pushes researchers to build models that yield good generalization performance and can be used in different conditions. In Tab. II, we also report average scores for mIoU and PQ across all four splits and both instance segmentation tasks, to show how different approaches deal with both tasks at the same time.
@@ V-C Ablation Studies
In the ablation study, we aim to validate our claim that the proposed skip connection scheme allows the CNN to exploit the underlying hierarchical relationship between the individual segmentation tasks, leading to better segmentation performance. We perform this experiment on the validation set. We compared with other skip connection schemes, and report results in Tab. IV. The simplest approach does not use skip connections (denoted as A in Tab. IV), which leads to poor performance. This is expected, since the decoders do not get any high-level information from features coming from earlier stages, which harms gradient flow. The second approach uses skip connections only from decoder to decoder, without any influence from the encoder (denoted as B in Tab. IV). This approach, despite exploiting the hierarchy among tasks with decoder-based skip connections, does not yield good performance. The third approach is the standard UNet-like skip connections from the encoder to all decoders (denoted as C in Tab. IV). In this case, all decoders obtain the same information from the encoder. This proves more effective than the decoder-only skip connections. This is probably due to the fact that restoring high-level information is important, and using decoder-only skip connections does not help gradient propagation, despite bringing valuable hierarchical information. Thus, our method using skip connections from the encoder and from decoder to decoder (denoted as D in Tab. IV) performs better than all others, as on one hand it benefits from the high-level features coming from the encoder, and on the other hand it exploits the underlying hierarchy among segmentation tasks. Interestingly, it also provides better results on semantic segmentation, despite there is no difference in the skip connection scheme for the semantic decoder, as the hierarchy affects only the instance decoders. This suggests that having information coming from the instance decoders indirectly affecting the encoder and the semantic decoder via backpropagation is useful for semantic segmentation performance.

### v2

@@ Abstract
Crop yield estimation is a relevant problem in agriculture, because an accurate yield estimate can support farmers’ decisions on harvesting or precision intervention. Robots can help to automate this process. To do so, they need to be able to perceive the surrounding environment to identify target objects such as trees and plants. In this paper, we introduce a novel approach to address the problem of hierarchical panoptic segmentation of apple orchards on 3D data from different sensors. Our approach is able to simultaneously provide semantic segmentation, instance segmentation of trunks and fruits, and instance segmentation of trees (a trunk with its fruits). This allows us to identify relevant information such as individual plants, fruits, and trunks, and capture the relationship among them, such as precisely estimate the number of fruits associated to each tree in an orchard. To efficiently evaluate our approach for hierarchical panoptic segmentation, we provide a dataset designed specifically for this task. Our dataset is recorded in Bonn, Germany, in a real apple orchard with a variety of sensors, spanning from a terrestrial laser scanner to a RGB-D camera mounted on different robots platforms. The experiments show that our approach surpasses state-of-the-art approaches in 3D panoptic segmentation in the agricultural domain, while also providing full hierarchical panoptic segmentation. Our dataset is publicly available at https://www.ipb.uni-bonn.de/data/hops/. The open-source implementation of our approach is available at https://github.com/PRBonn/hapt3D.
@@ I Introduction
Crop yield estimation is an important task in agriculture. An accurate yield estimate can help farmers making management decisions to increase crop production and optimize key factors such as harvest time and fertilizer use [16, 41, 44], as well as enable precision interventions [2]. Robots can automatize or support many of these interventions, but to do so reliably and robustly, they need to understand their surroundings through the interpretation of sensor data. Tasks such as segmenting, counting, and localizing fruits in orchards are crucial in horticulture, as they are required for replacing human fruit picking, which is an extremely labor-intensive process [3].
@@
Our experiments suggest that (i) our approach can jointly perform semantic, tree instance, and standard instance segmentation on real-world 3D data acquired with different sensors; and (ii) our skip connection scheme allows the CNN to exploit the underlying hierarchical relationship between the individual segmentation tasks, leading to better segmentation performance. Our claims are backed up by the paper and our experimental evaluation.
@@ II Related Work
Panoptic segmentation [14] unifies semantic and instance segmentation provides semantics of background classes, so-called “stuff”, and object-level instances of so-called “things” at the same time. While often images are used for panoptic segmentation in the agricultural domain [43], 3D point clouds generated by RGB-D cameras, LiDARs, or photogrammetric structure-from-motion approaches received increasing research interest [46] due to the capability to extract precise geometric information required for high-precision fruit grasping [22, 15], plant segmentation [30], and fine-grained trait estimation [26]. 3D panoptic segmentation has been actively investigated in the domain of autonomous driving [37, 28, 45], where most approaches use dedicated branches for semantic and instance segmentation using features of a shared encoder. For plant phenotyping, we can exploit the hierarchical structure of plants that can be decomposed in individual parts [11], such as plant, leaf and even more fine-grained leaf structure. This hierarchical structure has been exploited in the literature, leveraging skip connections between decoders for hierarchical bottom-up instance prediction tasks using images [34], which is the foundation of our approach to 3D panoptic segmentation.
In contrast to prior approaches for 3D panoptic segmentation, we also explicitly consider the hierarchical structure of the prediction task by means of our novel skip connections scheme, which allows us to predict a hierarchical semantic scene interpretation consisting of plant instances as well as fruit instances that can be used for plant monitoring and yield estimation in the agricultural domain. While our approach is applied to orchard scenes, the hierarchical panoptic segmentation framework is general and can be applied to other domains, such as body part segmentation [17].
While most agricultural datasets for semantic interpretation provide only RGB images [21], recently 3D datasets became available. In particular, BUP20 [38] provides an RGB-D dataset for instance segmentation of sweet peppers in a glasshouse [39]. For 3D plant phenotyping in real agricultural fields, BonnBeetClouds [25] provides annotated structure-from-motion point clouds consisting of plant and leaf instance of sugar beet plants. Other datasets provide organ-level annotations of plants using point clouds acquired with high-precision LiDAR scanners [36] or x-ray imaging [7] in the lab. The Crops3D dataset [46] provides organ-level annotations of different crop varieties for point cloud data acquired with a terrestrial laser scanner (TLS) in the field and single plant point clouds using structure-from-motion of RGB images or using structured light cameras.
@@ III-B Skip Connections
Skip connections are crucial to ensure feature reusability and address the gradient degradation problem of CNNs. They skip one or more layers, providing a direct gradient flow from the late stages of the decoder to the early stages of the encoder, which is usually hindered by the downsampling operation of the encoder. We aim to extend our previous work on hierarchical panoptic segmentation on RGB images [34] to the 3D point cloud domain, and thus we propose to adopt a hierarchical skip connections scheme to address the underlying hierarchy of segmentation tasks. We propose to connect different decoders directly as shown in Fig. 2, instead of encoder and decoders only.
@@
To better illustrate the information flow in our hierarchical panoptic segmentation architecture, consider the process as assembling segmentation in layers. The encoder provides a basic high-resolution spatial guide. The semantic decoder adds labels to it. The tree instance segmentation decoder builds upon this enriched map, identifying individual trees as landmarks. Finally, the instance segmentation decoder refines this by adding finer distinctions, such as fruits and trunks, using all previously gathered information. At each stage, skip connections act like reference notes, ensuring that each level benefits from both low-level detail and high-level context, progressively enriching the representation. This skip connection scheme allows us to exploit the hierarchy among tasks, enriching the features propagated to each decoder. In Sec. V-C, we provide extensive experiments on different skip connection schemes to show that this design choice yields superior performance.
@@ IV Our Dataset for
Our dataset, called HOPS, is composed of point clouds collected with three different sensors at Campus Klein-Altendorf near Bonn, Germany. First, we collected point clouds using a TLS placed at multiple locations in the orchard row. Our training and validation set entirely contains point clouds collected with this sensor. Furthermore, we use a few TLS point clouds in the test set, collected in a different year and orchard location compared to the training and validation set. Second, we use point clouds obtained with a bundle adjustment procedure [40] using as input images collected with a UAV equipped with a PhaseOne iXM-100 camera. We flew three missions on the same orchard with a camera angle of 45∘, 90∘, and 135∘ from the ground plane at a height of approximately 20 m. This setup allows us to obtain good coverage of the trees including the lower part of the canopy. The photogrammetric point clouds obtained in this way from the UAV are only included in the test set. Third, we use a RealSense d435i mounted on a mobile robot driving through the orchard rows to collect another part of our test set. Again, to obtain photogrammetric point clouds we use a bundle adjustment pipeline. Furthermore, we label a few point clouds from the Fuji-SfM dataset [9, 10] to obtain a fourth test set from a different setup. This dataset is collected in Agramunt in Catalunia, Spain with a Canon EOS 60D DSLR camera, followed by a bundle adjustment procedure to generate point clouds. In contrast to existing datasets in the agricultural domain [12] [33] [42], we specifically designed HOPS to have a remarkable domain shift between train and test sets, to more closely resemble real-world deployments. For this reason, we only included high-quality point clouds recorded with the TLS in the training set, while the test set includes a variety of sensors.
@@
To label the data, we split the aggregated point cloud of each orchard row in individual tiles with an average of 1 million points each. We manually label each tile in three stages using the online tool provided by segments.ai. One annotator labels each tile for semantic and two-level instance segmentation, ensuring consistency between instance levels. A second annotator then verifies label quality. Unlike image annotation, point cloud labeling requires frequent viewpoint changes, making it more time-consuming. The average effort consists of h per tile, totaling h for labeling and h for verification.
@@ V-A Experimental Setup
Metrics. For semantic segmentation, we compute the intersection over union (IoU) [8] over all five classes of our dataset, which we discussed in Sec. IV, and report the mean IoU (mIoU) in the following tables. For the tree instance segmentation, we compute the single-class panoptic quality over the “tree” class, that we report as PQ. For the standard instance segmentation, we compute the panoptic quality over all classes, and report the mean panoptic quality as PQ in the tables. We also report the overall mean panoptic quality (mPQ) as the average between PQ and PQ.
@@
| Approach | TLS | Param | s/it | |||
| mIoU | PQ | PQ | mPQ | |||
| MaskPLS [24] | 28.2 | 42.8 | - | 45.6 | 1.8M | 4.0 |
| MaskPLS [24] | - | - | 48.9 | |||
| ForestPS [23] | 58.9 | 62.2 | - | 32.3 | 1.5M | 1.8 |
| ForestPS [23] | - | - | 2.3 | |||
| Ours | 70.9 | 60.8 | 52.1 | 56.5 | 19.4M | 2.0 |
Training details and parameters. In all experiments, we use AdamW [18] with weight decay of with an initial learning rate of . We set the weights of the loss function to . We train our approach for 500 epochs. We set batch size of 1. We use voxel downsampling to . Due to practical constraints, the training set is limited to TLS-acquired data. However, to mitigate potential overfitting to TLS-specific characteristics and improve generalization to non-TLS domains, we apply extensive data augmentation techniques including scale, rotation around all axes, shear, color jittering, and elastic deformation [4]. We tuned all hyperparameters on the validation set.
@@ V-B Experiments on Hierarchical Panoptic Segmentation
While our PQ is inferior to ForestPS, our approach also yields good performance on tree instance segmentation, on which the dedicated training of ForestPS fails completely, achieving results below on four out of five splits. The first experiment evaluates the performance of our approach and its outcomes support the claim that our approach can jointly perform semantic, tree instance, and standard instance segmentation on real-world 3D data. For this experiment, we compare our approach to existing baselines for panoptic segmentation: MaskPLS [24] is a transformer-based approach that extends Mask2Former [5] to 3D point clouds, while ForestPS [23] is a convolutional neural network approach based on MinkUNet [6], originally designed for tree segmentation in a forest environment. The baselines are trained for only one of the two instance segmentation tasks at a time, as they do not support hierarchical multi-level instance segmentation. As shown in Tab. II for the test set, and Tab. III for the validation set, our approach outperforms the baselines on both splits in terms of mIoU and on most PQ, while ForestPS achieves the best PQ on most splits. However, our approach performs better than the baselines on mPQ.
We show qualitative results in Fig. 4. As explained in Sec. IV, only the validation set is recorded in the same time period and with the same sensor as the training set. The test sets, in contrast, are recorded either 2 years later also using a TLS, or with different sensors (UGV and UAV), or belong to entirely different datasets (SfM). This motivates the performance gap between the validation and the test set. We believe that this aspect pushes researchers to build models that yield good generalization performance and can be used in different conditions. In Tab. II, we also report average scores for mIoU and PQ across all four splits and both instance segmentation tasks, to show how different approaches deal with both tasks at the same time.
@@ V-C Ablation Studies
In the ablation study, we aim to validate our claim that the proposed skip connection scheme allows the CNN to exploit the underlying hierarchical relationship between the segmentation tasks, leading to better segmentation performance. We perform this experiment on the validation set. We compared to other skip connection schemes, and report results in Tab. IV. The simplest approach does not use skip connections (denoted as A in Tab. IV), which leads to poor performance. This is expected, since the decoders do not get any high-level information from features coming from earlier stages, which harms gradient flow. The second approach uses skip connections only from decoder to decoder, excluding the contributions from the encoder (denoted as B in Tab. IV). This approach, despite exploiting the hierarchy among tasks with decoder-based skip connections, does not yield good performance. The third approach is the standard UNet-like skip connections from the encoder to all decoders (denoted as C in Tab. IV). In this case, all decoders obtain the same information from the encoder. This proves more effective than the decoder-only skip connections. This is probably due to the fact that restoring high-level information is important, and using decoder-only skip connections does not help gradient propagation. Thus, our method using skip connections from the encoder and from decoder to decoder (denoted as D in Tab. IV) performs better than all others, as on one hand it benefits from the high-level features coming from the encoder, and on the other hand it exploits the underlying hierarchy among segmentation tasks. Interestingly, it also provides better results on semantic segmentation, despite there is no difference in the skip connection scheme for the semantic decoder, as the hierarchy affects only the instance decoders. This suggests that having information coming from the instance decoders indirectly affecting the encoder and the semantic decoder via backpropagation is useful for semantic segmentation performance.